The gnu Binary Utilities

Version 2.10.1

May 1993

Roland H. Pesch Jeffrey M. Osier Cygnus Support

Cygnus Support TEXinfo 1999-09-25.10

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 92, 93, 94, 95, 96, 97, 1998 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

Introduction

This brief manual contains preliminary documentation for the gnu binary utilities (collectively version 2.10.1): ar Create, modify, and extract from archives nm List symbols from object files objcopy Copy and translate object files objdump Display information from object files ranlib Generate index to archive contents readelf Display the contents of ELF format files. size List file section sizes and total size strings List printable strings from files strip Discard symbols c++filt Demangle encoded C++ symbols (on MS-DOS, this program is named cxxfilt)

addr2line

Convert addresses into file names and line numbers

nlmconv Convert object code into a Netware Loadable Module windres Manipulate Windows resources dlltool Create the files needed to build and use Dynamic Link Libraries 2 gnu Binary Utilities

1 ar

ar [-]p[mod [relpos] [count]] archive [member...]

     ar -M [ <mri-script ]

The gnu ar program creates, modifies, and extracts from archives. An archive is a single file holding a collection of other files in a structure that makes it possible to retrieve the original individual files (called members of the archive).

The original files’ contents, mode (permissions), timestamp, owner, and group are preserved in the archive, and can be restored on extraction.

gnu ar can maintain archives whose members have names of any length; however, depending on how ar is configured on your system, a limit on member-name length may be imposed for compatibility with archive formats maintained with other tools. If it exists, the limit is often 15 characters (typical of formats related to a.out) or 16 characters (typical of formats related to coff).

ar is considered a binary utility because archives of this sort are most often used as libraries holding commonly needed subroutines.

ar creates an index to the symbols defined in relocatable object modules in the archive when you specify the modifier ‘s’. Once created, this index is updated in the archive whenever ar makes a change to its contents (save for the ‘q’ update operation). An archive with such an index speeds up linking to the library, and allows routines in the library to call each other without regard to their placement in the archive.

You may use ‘nm -s’ or ‘nm --print-armap’ to list this index table. If an archive lacks the table, another form of ar called ranlib can be used to add just the table.

gnu ar is designed to be compatible with two different facilities. You can control its activity using command-line options, like the different varieties of ar on Unix systems; or, if you specify the single command-line option ‘-M’, you can control it with a script supplied via standard input, like the MRI “librarian” program.

1.1 Controlling ar on the command line

ar [-]p[mod [relpos] [count]] archive [member...] When you use ar in the Unix style, ar insists on at least two arguments to execute: one

keyletter specifying the operation (optionally accompanied by other keyletters specifying modifiers), and the archive name to act on.

Most operations can also accept further member arguments, specifying particular files to operate on.

gnu ar allows you to mix the operation code p and modifier flags mod in any order, within the first command-line argument.

If you wish, you may begin the first command-line argument with a dash. The p keyletter specifies what operation to execute; it may be any of the following, but

you must specify only one of them:

d Delete modules from the archive. Specify the names of modules to be deleted as member …; the archive is untouched if you specify no files to delete. If you specify the ‘v’ modifier, ar lists each module as it is deleted.

m Use this operation to move members in an archive.

The ordering of members in an archive can make a difference in how programs are linked using the library, if a symbol is defined in more than one member. If no modifiers are used with m, any members you name in the member arguments are moved to the end of the archive; you can use the ‘a’, ‘b’, or ‘i’ modifiers to move them to a specified place instead.

p Print the specified members of the archive, to the standard output file. If the ‘v’ modifier is specified, show the member name before copying its contents to standard output. If you specify no member arguments, all the files in the archive are printed.

q Quick append; Historically, add the files member … to the end of archive, without checking for replacement. The modifiers ‘a’, ‘b’, and ‘i’ do not affect this operation; new members are always placed at the end of the archive. The modifier ‘v’ makes ar list each file as it is appended. Since the point of this operation is speed, the archive’s symbol table index is not updated, even if it already existed; you can use ‘ar s’ or ranlib explicitly to update the symbol table index. However, too many different systems assume quick append rebuilds the index, so GNU ar implements q as a synonym for r.

r Insert the files member … into archive (with replacement). This operation differs from ‘q’ in that any previously existing members are deleted if their names match those being added. If one of the files named in member … does not exist, ar displays an error message, and leaves undisturbed any existing members of the archive matching that name.

By default, new members are added at the end of the file; but you may use one of the modifiers ‘a’, ‘b’, or ‘i’ to request placement relative to some existing member. The modifier ‘v’ used with this operation elicits a line of output for each file inserted, along with one of the letters ‘a’ or ‘r’ to indicate whether the file was appended (no old member deleted) or replaced.

t Display a table listing the contents of archive, or those of the files listed in member … that are present in the archive. Normally only the member name is shown; if you also want to see the modes (permissions), timestamp, owner, group, and size, you can request that by also specifying the ‘v’ modifier. If you do not specify a member, all files in the archive are listed. If there is more than one file with the same name (say, ‘fie’) in an archive (say ‘b.a’), ‘ar t b.a fie’ lists only the first instance; to see them all, you must ask for a complete listing—in our example, ‘ar t b.a’.

x Extract members (named member) from the archive. You can use the ‘v’ modifier with this operation, to request that ar list each name as it extracts it. If you do not specify a member, all files in the archive are extracted.

A number of modifiers (mod) may immediately follow the p keyletter, to specify variations on an operation’s behavior:

a Add new files after an existing member of the archive. If you use the modifier ‘a’, the name of an existing archive member must be present as the relpos argument, before the archive specification.

b Add new files before an existing member of the archive. If you use the modifier ‘b’, the name of an existing archive member must be present as the relpos argument, before the archive specification. (same as ‘i’).

c Create the archive. The specified archive is always created if it did not exist, when you request an update. But a warning is issued unless you specify in advance that you expect to create it, by using this modifier.

f Truncate names in the archive. gnu ar will normally permit file names of any length. This will cause it to create archives which are not compatible with the native ar program on some systems. If this is a concern, the ‘f’ modifier may be used to truncate file names when putting them in the archive.

i Insert new files before an existing member of the archive. If you use the modifier ‘i’, the name of an existing archive member must be present as the relpos argument, before the archive specification. (same as ‘b’).

l This modifier is accepted but not used.

N Uses the count parameter. This is used if there are multiple entries in the archive with the same name. Extract or delete instance count of the given name from the archive.

o Preserve the original dates of members when extracting them. If you do not specify this modifier, files extracted from the archive are stamped with the time of extraction.

P Use the full path name when matching names in the archive. gnu ar can not create an archive with a full path name (such archives are not POSIX complaint), but other archive creators can. This option will cause gnu ar to match file names using a complete path name, which can be convenient when extracting a single file from an archive created by another tool.

s Write an object-file index into the archive, or update an existing one, even if no other change is made to the archive. You may use this modifier flag either with any operation, or alone. Running ‘ar s’ on an archive is equivalent to running ‘ranlib’ on it.

S Do not generate an archive symbol table. This can speed up building a large library in several steps. The resulting archive can not be used with the linker. In order to build a symbol table, you must omit the ‘S’ modifier on the last execution of ‘ar’, or you must run ‘ranlib’ on the archive.

u Normally, ‘ar r’… inserts all files listed into the archive. If you would like to insert only those of the files you list that are newer than existing members of the same names, use this modifier. The ‘u’ modifier is allowed only for the operation ‘r’ (replace). In particular, the combination ‘qu’ is not allowed, since checking the timestamps would lose any speed advantage from the operation ‘q’.

v This modifier requests the verbose version of an operation. Many operations display additional information, such as filenames processed, when the modifier ‘v’ is appended.

V This modifier shows the version number of ar.

1.2 Controlling ar with a script

ar -M [ <script ]

If you use the single command-line option ‘-M’ with ar, you can control its operation with a rudimentary command language. This form of ar operates interactively if standard input is coming directly from a terminal. During interactive use, ar prompts for input (the prompt is ‘AR >’), and continues executing even after errors. If you redirect standard input to a script file, no prompts are issued, and ar abandons execution (with a nonzero exit code) on any error.

The ar command language is not designed to be equivalent to the command-line options; in fact, it provides somewhat less control over archives. The only purpose of the command language is to ease the transition to gnu ar for developers who already have scripts written for the MRI “librarian” program.

The syntax for the ar command language is straightforward:

• commands are recognized in upper or lower case; for example, LIST is the same as list. In the following descriptions, commands are shown in upper case for clarity.

• a single command may appear on each line; it is the first word on the line.

• empty lines are allowed, and have no effect.

• comments are allowed; text after either of the characters ‘*’ or ‘;’ is ignored.

• Whenever you use a list of names as part of the argument to an ar command, you can separate the individual names with either commas or blanks. Commas are shown in the explanations below, for clarity.

• ‘+’ is used as a line continuation character; if ‘+’ appears at the end of a line, the text on the following line is considered part of the current command.

Here are the commands you can use in ar scripts, or when using ar interactively. Three of them have special significance:

OPEN or CREATE specify a current archive, which is a temporary file required for most of the other commands.

SAVE commits the changes so far specified by the script. Prior to SAVE, commands affect only the temporary copy of the current archive.

ADDLIB archive ADDLIB archive (module, module, ... module)

Add all the contents of archive (or, if specified, each named module from archive) to the current archive. Requires prior use of OPEN or CREATE.

ADDMOD member, member, ... member Add each named member as a module in the current archive. Requires prior use of OPEN or CREATE.

CLEAR Discard the contents of the current archive, canceling the effect of any operations since the last SAVE. May be executed (with no effect) even if no current archive is specified.

CREATE archive Creates an archive, and makes it the current archive (required for many other commands). The new archive is created with a temporary name; it is not actually saved as archive until you use SAVE. You can overwrite existing archives; similarly, the contents of any existing file named archive will not be destroyed until SAVE.

DELETE module, module, ... module Delete each listed module from the current archive; equivalent to ‘ar -d archive module ... module’. Requires prior use of OPEN or CREATE.

DIRECTORY archive (module, ... module) DIRECTORY archive (module, ... module) outputfile

List each named module present in archive. The separate command VERBOSE specifies the form of the output: when verbose output is off, output is like that of ‘ar -t archive module...’. When verbose output is on, the listing is like ‘ar -tv archive module...’. Output normally goes to the standard output stream; however, if you specify outputfile as a final argument, ar directs the output to that file.

END Exit from ar, with a 0 exit code to indicate successful completion. This command does not save the output file; if you have changed the current archive since the last SAVE command, those changes are lost.

EXTRACT module, module, ... module Extract each named module from the current archive, writing them into the current directory as separate files. Equivalent to ‘ar -x archive module...’. Requires prior use of OPEN or CREATE.

LIST Display full contents of the current archive, in “verbose” style regardless of the state of VERBOSE. The effect is like ‘ar tv archive’. (This single command is a gnu ar enhancement, rather than present for MRI compatibility.) Requires prior use of OPEN or CREATE.

OPEN archive Opens an existing archive for use as the current archive (required for many other commands). Any changes as the result of subsequent commands will not actually affect archive until you next use SAVE.

REPLACE module, module, ... module In the current archive, replace each existing module (named in the REPLACE arguments) from files in the current working directory. To execute this command without errors, both the file, and the module in the current archive, must exist. Requires prior use of OPEN or CREATE.

VERBOSE Toggle an internal flag governing the output from DIRECTORY. When the flag is on, DIRECTORY output matches output from ‘ar -tv ’. . ..

SAVE Commit your changes to the current archive, and actually save it as a file with the name specified in the last CREATE or OPEN command. Requires prior use of OPEN or CREATE.

Chapter 2: ld 9

2 ld

The gnu linker ld is now described in a separate manual. See section “Overview” in Using LD: the gnu linker.

10 gnu Binary Utilities

3 nm

 nm [ -a | --debug-syms ]  [ -g | --extern-only ]
    [ -B ]  [ -C | --demangle ] [ -D | --dynamic ]
    [ -s | --print-armap ]  [ -A | -o | --print-file-name ]
    [ -n | -v | --numeric-sort ]  [ -p | --no-sort ]
    [ -r | --reverse-sort ]  [ --size-sort ] [ -u | --undefined-only ]
    [ -t radix | --radix=radix ] [ -P | --portability ]
    [ --target=bfdname ] [ -f format | --format=format ]
    [ --defined-only ] [-l | --line-numbers ]
    [ --no-demangle ] [ -V | --version ]  [ --help ]  [ objfile... ]

gnu nm lists the symbols from object files objfile. . .. If no object files are listed as arguments, nm assumes the file ‘a.out’.

For each symbol, nm shows:

• The symbol value, in the radix selected by options (see below), or hexadecimal by default.

• The symbol type. At least the following types are used; others are, as well, depending on the object file format. If lowercase, the symbol is local; if uppercase, the symbol is global (external).

A The symbol’s value is absolute, and will not be changed by further linking.

B The symbol is in the uninitialized data section (known as BSS).

C The symbol is common. Common symbols are uninitialized data. When linking, multiple common symbols may appear with the same name. If the symbol is defined anywhere, the common symbols are treated as undefined references. For more details on common symbols, see the discussion of –warn-common in section “Linker options” in The GNU linker.

D The symbol is in the initialized data section.

G The symbol is in an initialized data section for small objects. Some object file formats permit more efficient access to small data objects, such as a global int variable as opposed to a large global array.

I The symbol is an indirect reference to another symbol. This is a GNU extension to the a.out object file format which is rarely used.

N The symbol is a debugging symbol.

R The symbol is in a read only data section.

S The symbol is in an uninitialized data section for small objects.

T The symbol is in the text (code) section.

U The symbol is undefined.

V The symbol is a weak object. When a weak defined symbol is linked with a normal defined symbol, the normal defined symbol is used with no error. When a weak undefined symbol is linked and the symbol is not defined, the value of the weak symbol becomes zero with no error.

W The symbol is a weak symbol that has not been specifically tagged as a weak object symbol. When a weak defined symbol is linked with a normal defined symbol, the normal defined symbol is used with no error. When a weak undefined symbol is linked and the symbol is not defined, the value of the weak symbol becomes zero with no error.

- The symbol is a stabs symbol in an a.out object file. In this case, the next values printed are the stabs other field, the stabs desc field, and the stab type. Stabs symbols are used to hold debugging information; for more information, see section “Stabs Overview” in The “stabs” debug format.

? The symbol type is unknown, or object file format specific.

• The symbol name.

The long and short forms of options, shown here as alternatives, are equivalent.

-A
-o
--print-file-name

Precede each symbol by the name of the input file (or archive member) in which it was found, rather than identifying the input file once only, before all of its symbols.

-a
--debug-syms

Display all symbols, even debugger-only symbols; normally these are not listed.

-B The same as ‘--format=bsd’ (for compatibility with the MIPS nm).

-C
--demangle

Decode (demangle) low-level symbol names into user-level names. Besides removing any initial underscore prepended by the system, this makes C++ function names readable. See Chapter 10 [c++filt], page 35, for more information on demangling.

--no-demangle

Do not demangle low-level symbol names. This is the default.

-D
--dynamic

Display the dynamic symbols rather than the normal symbols. This is only meaningful for dynamic objects, such as certain types of shared libraries.

-f format

--format=format

Use the output format format, which can be bsd, sysv, or posix. The default is bsd. Only the first character of format is significant; it can be either upper or lower case.

-g
--extern-only

Display only external symbols.

-l
--line-numbers

For each symbol, use debugging information to try to find a filename and line number. For a defined symbol, look for the line number of the address of the symbol. For an undefined symbol, look for the line number of a relocation entry which refers to the symbol. If line number information can be found, print it after the other symbol information.

-n
-v
--numeric-sort

Sort symbols numerically by their addresses, rather than alphabetically by their names.

-p
--no-sort

Do not bother to sort the symbols in any order; print them in the order encountered.

-P
--portability

Use the POSIX.2 standard output format instead of the default format. Equivalent to ‘-f posix’.

-s
--print-armap

When listing symbols from archive members, include the index: a mapping (stored in the archive by ar or ranlib) of which modules contain definitions for which names.

-r
--reverse-sort

Reverse the order of the sort (whether numeric or alphabetic); let the last come first.

--size-sort

Sort symbols by size. The size is computed as the difference between the value of the symbol and the value of the symbol with the next higher value. The size of the symbol is printed, rather than the value.

-t radix

--radix=radix

Use radix as the radix for printing the symbol values. It must be ‘d’ for decimal, ‘o’ for octal, or ‘x’ for hexadecimal.

--target=bfdname

Specify an object code format other than your system’s default format. See Section 16.1 [Target Selection], page 51, for more information.

-u
--undefined-only

Display only undefined symbols (those external to each object file).

--defined-only

Display only defined symbols for each object file.

-V
--version

Show the version number of nm and exit.

--help Show a summary of the options to nm and exit.

4 objcopy

objcopy [ -F bfdname | --target=bfdname ]
[ -I bfdname | --input-target=bfdname ]
[ -O bfdname | --output-target=bfdname ]
[ -S | --strip-all ]  [ -g | --strip-debug ]

[ -K symbolname | --keep-symbol=symbolname ]

[ -N symbolname | --strip-symbol=symbolname ]
[ -L symbolname | --localize-symbol=symbolname ]
[ -W symbolname | --weaken-symbol=symbolname ]
[ -x | --discard-all ]  [ -X | --discard-locals ]
[ -b byte | --byte=byte ]

[ -i interleave | --interleave=interleave ] [ -j sectionname | --only-section=sectionname ]

[ -R sectionname | --remove-section=sectionname ]
[ -p | --preserve-dates ] [ --debugging ]
[ --gap-fill=val ] [ --pad-to=address ]
[ --set-start=val ] [ --adjust-start=incr ]
[ --change-addresses=incr ]
[ --change-section-address section{=,+,-}val ]
[ --change-section-lma section{=,+,-}val ]
[ --change-section-vma section{=,+,-}val ]
[ --change-warnings ] [ --no-change-warnings ]
[ --set-section-flags section=flags ]

[ --add-section sectionname=filename ]

[ --change-leading-char ] [ --remove-leading-char ]
[ --redefine-sym old=new ] [ --weaken ]
[ -v | --verbose ] [ -V | --version ]  [ --help ]

infile [outfile]

The gnu objcopy utility copies the contents of an object file to another. objcopy uses the gnu bfd Library to read and write the object files. It can write the destination object file in a format different from that of the source object file. The exact behavior of objcopy is controlled by command-line options.

objcopy creates temporary files to do its translations and deletes them afterward. objcopy uses bfd to do all its translation work; it has access to all the formats described in bfd and thus is able to recognize most formats without being told explicitly. See section “BFD” in Using LD.

objcopy can be used to generate S-records by using an output target of ‘srec’ (e.g., use

‘-O srec’).

objcopy can be used to generate a raw binary file by using an output target of ‘binary’ (e.g., use ‘-O binary’). When objcopy generates a raw binary file, it will essentially produce a memory dump of the contents of the input object file. All symbols and relocation information will be discarded. The memory dump will start at the load address of the lowest section copied into the output file.

When generating an S-record or a raw binary file, it may be helpful to use ‘-S’ to remove sections containing debugging information. In some cases ‘-R’ will be useful to remove sections which contain information that is not needed by the binary file. infile outfile The input and output files, respectively. If you do not specify outfile, objcopy

creates a temporary file and destructively renames the result with the name of infile.

-I bfdname

--input-target=bfdname

Consider the source file’s object format to be bfdname, rather than attempting to deduce it. See Section 16.1 [Target Selection], page 51, for more information.

-O bfdname

--output-target=bfdname

Write the output file using the object format bfdname. See Section 16.1 [Target Selection], page 51, for more information.

-F bfdname

--target=bfdname

Use bfdname as the object format for both the input and the output file; i.e., simply transfer data from source to destination with no translation. See Section 16.1 [Target Selection], page 51, for more information.

-j sectionname

--only-section=sectionname

Copy only the named section from the input file to the output file. This option may be given more than once. Note that using this option inappropriately may make the output file unusable.

-R sectionname

--remove-section=sectionname

Remove any section named sectionname from the output file. This option may be given more than once. Note that using this option inappropriately may make the output file unusable.

-S
--strip-all

Do not copy relocation and symbol information from the source file.

-g
--strip-debug

Do not copy debugging symbols from the source file.

--strip-unneeded

Strip all symbols that are not needed for relocation processing.

-K symbolname

--keep-symbol=symbolname

Copy only symbol symbolname from the source file. This option may be given more than once.

-N symbolname

--strip-symbol=symbolname

Do not copy symbol symbolname from the source file. This option may be given more than once.

-L symbolname

--localize-symbol=symbolname

Make symbol symbolname local to the file, so that it is not visible externally. This option may be given more than once.

-W symbolname

--weaken-symbol=symbolname

Make symbol symbolname weak. This option may be given more than once.

-x
--discard-all

Do not copy non-global symbols from the source file.

-X
--discard-locals

Do not copy compiler-generated local symbols. (These usually start with ‘L’ or ‘.’.)

-b byte

--byte=byte

Keep only every byteth byte of the input file (header data is not affected). byte can be in the range from 0 to interleave-1, where interleave is given by the ‘-i’ or ‘--interleave’ option, or the default of 4. This option is useful for creating files to program rom. It is typically used with an srec output target.

-i interleave

--interleave=interleave

Only copy one out of every interleave bytes. Select which byte to copy with the -b or ‘--byte’ option. The default is 4. objcopy ignores this option if you do not specify either ‘-b’ or ‘--byte’.

-p
--preserve-dates

Set the access and modification dates of the output file to be the same as those of the input file.

--debugging

Convert debugging information, if possible. This is not the default because only certain debugging formats are supported, and the conversion process can be time consuming.

--gap-fill val

Fill gaps between sections with val. This operation applies to the load address (LMA) of the sections. It is done by increasing the size of the section with the lower address, and filling in the extra space created with val.

--pad-to address

Pad the output file up to the load address address. This is done by increasing the size of the last section. The extra space is filled in with the value specified by ‘--gap-fill’ (default zero).

--set-start val

Set the start address of the new file to val. Not all object file formats support setting the start address.

--change-start incr
--adjust-start incr

Change the start address by adding incr. Not all object file formats support setting the start address.

--change-addresses incr
--adjust-vma incr

Change the VMA and LMA addresses of all sections, as well as the start address, by adding incr. Some object file formats do not permit section addresses to be changed arbitrarily. Note that this does not relocate the sections; if the program expects sections to be loaded at a certain address, and this option is used to change the sections such that they are loaded at a different address, the program may fail.

--change-section-address section{=,+,-}val
--adjust-section-vma section{=,+,-}val

Set or change both the VMA address and the LMA address of the named section. If ‘=’ is used, the section address is set to val. Otherwise, val is added to or subtracted from the section address. See the comments under ‘--change-addresses’, above. If section does not exist in the input file, a warning will be issued, unless ‘--no-change-warnings’ is used.

--change-section-lma section{=,+,-}val

Set or change the LMA address of the named section. The LMA address is the address where the section will be loaded into memory at program load time. Normally this is the same as the VMA address, which is the address of the section at program run time, but on some systems, especially those where a program is held in ROM, the two can be different. If ‘=’ is used, the section address is set to val. Otherwise, val is added to or subtracted from the section address. See the comments under ‘--change-addresses’, above. If section does not exist in the input file, a warning will be issued, unless

‘--no-change-warnings’ is used.
--change-section-vma section{=,+,-}val

Set or change the VMA address of the named section. The VMA address is the address where the section will be located once the program has started executing. Normally this is the same as the LMA address, which is the address where the section will be loaded into memory, but on some systems, especially those where a program is held in ROM, the two can be different. If ‘=’ is used, the section address is set to val. Otherwise, val is added to or subtracted from the section address. See the comments under ‘--change-addresses’, above. If section does not exist in the input file, a warning will be issued, unless

‘--no-change-warnings’ is used.

--change-warnings
--adjust-warnings
If ‘--change-section-address’ or ‘--change-section-lma’ or

‘--change-section-vma’ is used, and the named section does not exist, issue a warning. This is the default.

--no-change-warnings
--no-adjust-warnings
Do not issue a warning if ‘--change-section-address’ or
‘--adjust-section-lma’ or ‘--adjust-section-vma’ is used, even if

the named section does not exist.

--set-section-flags section=flags

Set the flags for the named section. The flags argument is a comma separated string of flag names. The recognized names are ‘alloc’, ‘contents’, ‘load’,

‘noload’, ‘readonly’, ‘code’, ‘data’, ‘rom’, ‘share’, and ‘debug’. You can set

the ‘contents’ flag for a section which does not have contents, but it is not meaningful to clear the ‘contents’ flag of a section which does have contents– just remove the section instead. Not all flags are meaningful for all object file formats.

--add-section sectionname=filename Add a new section named sectionname while copying the file. The contents of the new section are taken from the file filename. The size of the section will be the size of the file. This option only works on file formats which can support sections with arbitrary names.

--change-leading-char

Some object file formats use special characters at the start of symbols. The most common such character is underscore, which compilers often add before every symbol. This option tells objcopy to change the leading character of every symbol when it converts between object file formats. If the object file formats use the same leading character, this option has no effect. Otherwise, it will add a character, or remove a character, or change a character, as appropriate.

--remove-leading-char

If the first character of a global symbol is a special symbol leading character used by the object file format, remove the character. The most common symbol leading character is underscore. This option will remove a leading underscore from all global symbols. This can be useful if you want to link together objects of different file formats with different conventions for symbol names. This is different from --change-leading-char because it always changes the symbol name when appropriate, regardless of the object file format of the output file.

--redefine-sym old=new

Change the name of a symbol old, to new. This can be useful when one is trying link two things together for which you have no source, and there are name collisions.

--weaken Change all global symbols in the file to be weak. This can be useful when building an object which will be linked against other objects using the -R option to the linker. This option is only effective when using an object file format which supports weak symbols.

-V
--version

Show the version number of objcopy.

-v
--verbose

Verbose output: list all object files modified. In the case of archives, ‘objcopy -V’ lists all members of the archive.

--help Show a summary of the options to objcopy.

5 objdump

objdump [ -a | --archive-headers ]
[ -b bfdname | --target=bfdname ]
[ -C | --demangle ]
[ -d | --disassemble ]
[ -D | --disassemble-all ]
[ -z | --disassemble-zeroes ]
[ -EB | -EL | --endian={big | little } ]
[ -f | --file-headers ]
[ --file-start-context ]
[ -g | --debugging ]
[ -h | --section-headers | --headers ]
[ -i | --info ]
[ -j section | --section=section ]
[ -l | --line-numbers ]
[ -S | --source ]
[ -m machine | --architecture=machine ]
[ -M options | --disassembler-options=options]
[ -p | --private-headers ]
[ -r | --reloc ]
[ -R | --dynamic-reloc ]
[ -s | --full-contents ]
[ -G | --stabs ]
[ -t | --syms ]
[ -T | --dynamic-syms ]
[ -x | --all-headers ]
[ -w | --wide ]
[ --start-address=address ]
[ --stop-address=address ]
[ --prefix-addresses]
[ --[no-]show-raw-insn ]
[ --adjust-vma=offset ]
[ -V | --version ]
[ -H | --help ]

objfile...

objdump displays information about one or more object files. The options control what particular information to display. This information is mostly useful to programmers who are working on the compilation tools, as opposed to programmers who just want their program to compile and work.

objfile… are the object files to be examined. When you specify archives, objdump shows information on each of the member object files.

The long and short forms of options, shown here as alternatives, are equivalent. At least

one option from the list ‘-a,-d,-D,-f,-g,-G,-h,-H,-p,-r,-R,-S,-t,-T,-V,-x’ must be

given.

-a
--archive-header

If any of the objfile files are archives, display the archive header information (in a format similar to ‘ls -l’). Besides the information you could list with ‘ar tv’, ‘objdump -a’ shows the object file format of each archive member.

--adjust-vma=offset

When dumping information, first add offset to all the section addresses. This is useful if the section addresses do not correspond to the symbol table, which can happen when putting sections at particular addresses when using a format which can not represent section addresses, such as a.out.

-b bfdname

--target=bfdname

Specify that the object-code format for the object files is bfdname. This option may not be necessary; objdump can automatically recognize many formats. For example,

     objdump -b oasys -m vax -h fu.o

displays summary information from the section headers (‘-h’) of ‘fu.o’, which is explicitly identified (‘-m’) as a VAX object file in the format produced by Oasys compilers. You can list the formats available with the ‘-i’ option. See Section 16.1 [Target Selection], page 51, for more information.

-C
--demangle

Decode (demangle) low-level symbol names into user-level names. Besides removing any initial underscore prepended by the system, this makes C++ function names readable. See Chapter 10 [c++filt], page 35, for more information on demangling.

-G
--debugging

Display debugging information. This attempts to parse debugging information stored in the file and print it out using a C like syntax. Only certain types of debugging information have been implemented.

-d
--disassemble

Display the assembler mnemonics for the machine instructions from objfile. This option only disassembles those sections which are expected to contain instructions.

-D
--disassemble-all

Like ‘-d’, but disassemble the contents of all sections, not just those expected to contain instructions.

--prefix-addresses

When disassembling, print the complete address on each line. This is the older disassembly format.

--disassemble-zeroes

Normally the disassembly output will skip blocks of zeroes. This option directs the disassembler to disassemble those blocks, just like any other data.

-EB
-EL
--endian={big|little}

Specify the endianness of the object files. This only affects disassembly. This can be useful when disassembling a file format which does not describe endianness information, such as S-records.

-f
--file-header

Display summary information from the overall header of each of the objfile files.

--file-start-context

Specify that when displaying interlisted source code/disassembly (assumes ’-S’) from a file that has not yet been displayed, extend the context to the start of the file.

-h
--section-header

--header Display summary information from the section headers of the object file. File segments may be relocated to nonstandard addresses, for example by using the ‘-Ttext’, ‘-Tdata’, or ‘-Tbss’ options to ld. However, some object file formats, such as a.out, do not store the starting address of the file segments. In those situations, although ld relocates the sections correctly, using ‘objdump -h’ to list the file section headers cannot show the correct addresses. Instead, it shows the usual addresses, which are implicit for the target.

--help Print a summary of the options to objdump and exit.

-i

--info Display a list showing all architectures and object formats available for specification with ‘-b’ or ‘-m’.

-j name

--section=name

Display information only for section name.

-l
--line-numbers

Label the display (using debugging information) with the filename and source line numbers corresponding to the object code or relocs shown. Only useful with ‘-d’, ‘-D’, or ‘-r’.

-m machine

--architecture=machine

Specify the architecture to use when disassembling object files. This can be useful when disassembling object files which do not describe architecture information, such as S-records. You can list the available architectures with the ‘-i’ option.

-M options

--disassembler-options=options

Pass target specific information to the disassembler. Only supported on some targets. If the target is an ARM architecture then this switch can be used to select which register name set is used during disassembler. Specifying ‘-M reg-name-std’ (the default) will select the register names as used in ARM’s instruction set documentation, but with register 13 called ’sp’, register 14 called ’lr’ and register 15 called ’pc’. Specifying ‘-M reg-names-apcs’ will select the name set used by the ARM Procedure Call Standard, whilst specifying ‘-M reg-names-raw’ will just use ‘r’ followed by the register number. There are also two variants on the APCS register naming scheme enabled

by ‘-M reg-names-atpcs’ and ‘-M reg-names-special-atpcs’ which use the

ARM/Thumb Procedure Call Standard naming conventions. (Eiuther with the normal register name sor the special register names). This option can also be used for ARM architectures to force the disassembler to interpret all instructions as THUMB instructions by using the switch

‘--disassembler-options=force-thumb’. This can be useful when attempt-

ing to disassemble thumb code produced by other compilers.

-p
--private-headers

Print information that is specific to the object file format. The exact information printed depends upon the object file format. For some object file formats, no additional information is printed.

-r

--reloc Print the relocation entries of the file. If used with ‘-d’ or ‘-D’, the relocations are printed interspersed with the disassembly.

-R
--dynamic-reloc

Print the dynamic relocation entries of the file. This is only meaningful for dynamic objects, such as certain types of shared libraries.

-s
--full-contents

Display the full contents of any sections requested.

-S

--source Display source code intermixed with disassembly, if possible. Implies ‘-d’.

--show-raw-insn

When disassembling instructions, print the instruction in hex as well as in symbolic form. This is the default except when --prefix-addresses is used.

--no-show-raw-insn

When disassembling instructions, do not print the instruction bytes. This is the default when --prefix-addresses is used.

-G

--stabs Display the full contents of any sections requested. Display the contents of the .stab and .stab.index and .stab.excl sections from an ELF file. This is only useful on systems (such as Solaris 2.0) in which .stab debugging symbol-table entries are carried in an ELF section. In most other file formats, debugging symbol-table entries are interleaved with linkage symbols, and are visible in the ‘--syms’ output. For more information on stabs symbols, see section “Stabs Overview” in The “stabs” debug format.

--start-address=address

Start displaying data at the specified address. This affects the output of the -d, -r and -s options.

--stop-address=address

Stop displaying data at the specified address. This affects the output of the -d, -r and -s options.

-t

--syms Print the symbol table entries of the file. This is similar to the information provided by the ‘nm’ program.

-T
--dynamic-syms

Print the dynamic symbol table entries of the file. This is only meaningful for dynamic objects, such as certain types of shared libraries. This is similar to the information provided by the ‘nm’ program when given the ‘-D’ (‘--dynamic’) option.

--version

Print the version number of objdump and exit.

-x
--all-header

Display all available header information, including the symbol table and relocation entries. Using ‘-x’ is equivalent to specifying all of ‘-a -f -h -r -t’.

-w

--wide Format some lines for output devices that have more than 80 columns.

26 gnu Binary Utilities

6 ranlib

     ranlib [-vV] archive

ranlib generates an index to the contents of an archive and stores it in the archive. The index lists each symbol defined by a member of an archive that is a relocatable object file.

You may use ‘nm -s’ or ‘nm --print-armap’ to list this index.

An archive with such an index speeds up linking to the library and allows routines in the library to call each other without regard to their placement in the archive.

The gnu ranlib program is another form of gnu ar; running ranlib is completely equivalent to executing ‘ar -s’. See Chapter 1 [ar], page 3.

-v
-V
--version

Show the version number of ranlib.

28 gnu Binary Utilities

7 size

size [ -A | -B | --format=compatibility ]
[ --help ]  [ -d | -o | -x | --radix=number ]
[ --target=bfdname ]  [ -V | --version ]

[ objfile... ]

The gnu size utility lists the section sizes—and the total size—for each of the object or archive files objfile in its argument list. By default, one line of output is generated for each object file or each module in an archive.

objfile… are the object files to be examined. If none are specified, the file a.out will be used.

The command line options have the following meanings:

-A
-B

--format=compatibility Using one of these options, you can choose whether the output from gnu size resembles output from System V size (using ‘-A’, or ‘--format=sysv’), or Berkeley size (using ‘-B’, or ‘--format=berkeley’). The default is the oneline format similar to Berkeley’s.

Here is an example of the Berkeley (default) format of output from size:

$ size --format=Berkeley ranlib size
text    data    bss     dec     hex     filename
294880  81920   11592   388392  5ed28   ranlib
294880  81920   11888   388688  5ee50   size

This is the same data, but displayed closer to System V conventions:

$ size --format=SysV ranlib size
ranlib  :
section         size         addr
.text         294880         8192
.data          81920       303104
.bss           11592       385024
Total         388392

size  :
section         size         addr
.text         294880         8192
.data          81920       303104
.bss           11888       385024
Total         388688

--help Show a summary of acceptable arguments and options.

-d
-o
-x
--radix=number

Using one of these options, you can control whether the size of each section is given in decimal (‘-d’, or ‘--radix=10’); octal (‘-o’, or ‘--radix=8’); or hexadecimal (‘-x’, or ‘--radix=16’). In ‘--radix=number’, only the three values (8, 10, 16) are supported. The total size is always given in two radices; decimal and hexadecimal for ‘-d’ or ‘-x’ output, or octal and hexadecimal if you’re using ‘-o’.

--target=bfdname

Specify that the object-code format for objfile is bfdname. This option may not be necessary; size can automatically recognize many formats. See Section 16.1 [Target Selection], page 51, for more information.

-V
--version

Display the version number of size.

8 strings

     strings [-afov] [-min-len] [-n min-len] [-t radix] [-]
             [--all] [--print-file-name] [--bytes=min-len]
             [--radix=radix] [--target=bfdname]
             [--help] [--version] file...

For each file given, gnu strings prints the printable character sequences that are at least 4 characters long (or the number given with the options below) and are followed by an unprintable character. By default, it only prints the strings from the initialized and loaded sections of object files; for other types of files, it prints the strings from the whole file.

strings is mainly useful for determining the contents of non-text files.

-a
--all

- Do not scan only the initialized and loaded sections of object files; scan the whole files.

-f
--print-file-name

Print the name of the file before each string.

--help Print a summary of the program usage on the standard output and exit.

-min-len -n min-len

--bytes=min-len

Print sequences of characters that are at least min-len characters long, instead of the default 4.

-o Like ‘-t o’. Some other versions of strings have ‘-o’ act like ‘-t d’ instead. Since we can not be compatible with both ways, we simply chose one.

-t radix

--radix=radix

Print the offset within the file before each string. The single character argument specifies the radix of the offset—‘o’ for octal, ‘x’ for hexadecimal, or ‘d’ for decimal.

--target=bfdname

Specify an object code format other than your system’s default format. See Section 16.1 [Target Selection], page 51, for more information.

-v
--version

Print the program version number on the standard output and exit.

32 gnu Binary Utilities

9 strip

     strip [ -F bfdname | --target=bfdname ]
           [ -I bfdname | --input-target=bfdname ]
           [ -O bfdname | --output-target=bfdname ]
           [ -s | --strip-all ] [ -S | -g | --strip-debug ]

[ -K symbolname | --keep-symbol=symbolname ]

           [ -N symbolname | --strip-symbol=symbolname ]
           [ -x | --discard-all ] [ -X | --discard-locals ]
           [ -R sectionname | --remove-section=sectionname ]
           [ -o file ] [ -p | --preserve-dates ]
           [ -v | --verbose ]  [ -V | --version ]  [ --help ]

objfile...

gnu strip discards all symbols from object files objfile. The list of object files may include archives. At least one object file must be given.

strip modifies the files named in its argument, rather than writing modified copies under different names.

-F bfdname

--target=bfdname

Treat the original objfile as a file with the object code format bfdname, and rewrite it in the same format. See Section 16.1 [Target Selection], page 51, for more information.

--help Show a summary of the options to strip and exit.

-I bfdname

--input-target=bfdname

Treat the original objfile as a file with the object code format bfdname. See Section 16.1 [Target Selection], page 51, for more information.

-O bfdname

--output-target=bfdname

Replace objfile with a file in the output format bfdname. See Section 16.1 [Target Selection], page 51, for more information.

-R sectionname

--remove-section=sectionname

Remove any section named sectionname from the output file. This option may be given more than once. Note that using this option inappropriately may make the output file unusable.

-s
--strip-all

Remove all symbols.

-g
-S
--strip-debug

Remove debugging symbols only.

--strip-unneeded

Remove all symbols that are not needed for relocation processing.

-K symbolname

--keep-symbol=symbolname

Keep only symbol symbolname from the source file. This option may be given more than once.

-N symbolname

--strip-symbol=symbolname

Remove symbol symbolname from the source file. This option may be given more than once, and may be combined with strip options other than -K.

-o file Put the stripped output in file, rather than replacing the existing file. When this argument is used, only one objfile argument may be specified.

-p
--preserve-dates

Preserve the access and modification dates of the file.

-x
--discard-all

Remove non-global symbols.

-X
--discard-locals

Remove compiler-generated local symbols. (These usually start with ‘L’ or ‘.’.)

-V
--version

Show the version number for strip.

-v
--verbose

Verbose output: list all object files modified. In the case of archives, ‘strip -v’ lists all members of the archive.

10 c++filt

     c++filt [ -_| --strip-underscores ]
             [ -j | --java ]
     [ -n | --no-strip-underscores ]
             [ -s format | --format=format ]
             [ --help ]  [ --version ]  [ symbol... ]

The C++ and Java languages provides function overloading, which means that you can write many functions with the same name (providing each takes parameters of different types). All C++ and Java function names are encoded into a low-level assembly label (this process is known as mangling). The c++filt^{1} program does the inverse mapping: it decodes (demangles) low-level names into user-level names so that the linker can keep these overloaded functions from clashing.

Every alphanumeric word (consisting of letters, digits, underscores, dollars, or periods) seen in the input is a potential label. If the label decodes into a C++ name, the C++ name replaces the low-level name in the output.

You can use c++filt to decipher individual symbols:

     c++filt symbol

If no symbol arguments are given, c++filt reads symbol names from the standard input and writes the demangled names to the standard output. All results are printed on the standard output.

-_
--strip-underscores

On some systems, both the C and C++ compilers put an underscore in front of every name. For example, the C name foo gets the low-level name _foo. This option removes the initial underscore. Whether c++filt removes the underscore by default is target dependent.

-j

--java Prints demangled names using Java syntax. The default is to use C++ syntax.

-n
--no-strip-underscores

Do not remove the initial underscore.

-s format

--format=format

gnu nm can decode three different methods of mangling, used by different C++ compilers. The argument to this option selects which method it uses:

gnu the one used by the gnu compiler (the default method)

lucid the one used by the Lucid compiler

arm the one specified by the C++ Annotated Reference Manual

hp the one used by the HP compiler

1 MS-DOS does not allow + characters in file names, so on MS-DOS this program is named cxxfilt.

edg the one used by the EDG compiler

--help Print a summary of the options to c++filt and exit.

--version

Print the version number of c++filt and exit.

Warning: c++filt is a new utility, and the details of its user interface are subject to change in future releases. In particular, a command-line option may be required in the the future to decode a name passed as an argument on the command line; in other words,

     c++filt symbol

may in a future release become c++filt option symbol

11 addr2line

     addr2line [ -b bfdname | --target=bfdname ]
               [ -C | --demangle ]

[ -e filename | --exe=filename ]

               [ -f | --functions ] [ -s | --basename ]
               [ -H | --help ] [ -V | --version ]
               [ addr addr ... ]

addr2line translates program addresses into file names and line numbers. Given an address and an executable, it uses the debugging information in the executable to figure out which file name and line number are associated with a given address.

The executable to use is specified with the -e option. The default is the file ‘a.out’.

addr2line has two modes of operation.

In the first, hexadecimal addresses are specified on the command line, and addr2line displays the file name and line number for each address.

In the second, addr2line reads hexadecimal addresses from standard input, and prints the file name and line number for each address on standard output. In this mode, addr2line may be used in a pipe to convert dynamically chosen addresses.

The format of the output is ‘FILENAME:LINENO’. The file name and line number for each address is printed on a separate line. If the -f option is used, then each ‘FILENAME:LINENO’ line is preceded by a ‘FUNCTIONNAME’ line which is the name of the function containing the address.

If the file name or function name can not be determined, addr2line will print two question marks in their place. If the line number can not be determined, addr2line will print 0.

The long and short forms of options, shown here as alternatives, are equivalent.

-b bfdname

--target=bfdname

Specify that the object-code format for the object files is bfdname.

-C
--demangle

Decode (demangle) low-level symbol names into user-level names. Besides removing any initial underscore prepended by the system, this makes C++ function names readable. See Chapter 10 [c++filt], page 35, for more information on demangling.

-e filename --exe=filename

Specify the name of the executable for which addresses should be translated. The default file is ‘a.out’.

-f
--functions

Display function names as well as file and line number information.

-s
--basenames

Display only the base of each file name.

12 nlmconv

nlmconv converts a relocatable object file into a NetWare Loadable Module. Warning: nlmconv is not always built as part of the binary utilities, since it is only useful for NLM targets.

     nlmconv [ -I bfdname | --input-target=bfdname ]
             [ -O bfdname | --output-target=bfdname ]
             [ -T headerfile | --header-file=headerfile ]
             [ -d | --debug]  [ -l linker | --linker=linker ]
             [ -h | --help ]  [ -V | --version ]

infile outfile

nlmconv converts the relocatable ‘i386’ object file infile into the NetWare Loadable Module outfile, optionally reading headerfile for NLM header information. For instructions on writing the NLM command file language used in header files, see the ‘linkers’ section, ‘NLMLINK’ in particular, of the NLM Development and Tools Overview, which is part of the NLM Software Developer’s Kit (“NLM SDK”), available from Novell, Inc. nlmconv uses the gnu Binary File Descriptor library to read infile; see section “BFD” in Using LD, for more information.

nlmconv can perform a link step. In other words, you can list more than one object file for input if you list them in the definitions file (rather than simply specifying one input file on the command line). In this case, nlmconv calls the linker for you.

-I bfdname

--input-target=bfdname

Object format of the input file. nlmconv can usually determine the format of a given file (so no default is necessary). See Section 16.1 [Target Selection], page 51, for more information.

-O bfdname

--output-target=bfdname

Object format of the output file. nlmconv infers the output format based on the input format, e.g. for a ‘i386’ input file the output format is ‘nlm32-i386’. See Section 16.1 [Target Selection], page 51, for more information.

-T headerfile

--header-file=headerfile

Reads headerfile for NLM header information. For instructions on writing the NLM command file language used in header files, see see the ‘linkers’ section, of the NLM Development and Tools Overview, which is part of the NLM Software Developer’s Kit, available from Novell, Inc.

-d

--debug Displays (on standard error) the linker command line used by nlmconv.

-l linker

--linker=linker

Use linker for any linking. linker can be an absolute or a relative pathname.

-h

--help Prints a usage summary.

-V
--version

Prints the version number for nlmconv.

13 windres

windres may be used to manipulate Windows resources.

Warning: windres is not always built as part of the binary utilities, since it is only useful for Windows targets.

     windres [options] [input-file] [output-file]

windres reads resources from an input file and copies them into an output file. Either file may be in one of three formats:

rc A text format read by the Resource Compiler.

res A binary format generated by the Resource Compiler.

coff A COFF object or executable.

The exact description of these different formats is available in documentation from Microsoft.

When windres converts from the rc format to the res format, it is acting like the Windows Resource Compiler. When windres converts from the res format to the coff format, it is acting like the Windows CVTRES program.

When windres generates an rc file, the output is similar but not identical to the format expected for the input. When an input rc file refers to an external filename, an output rc file will instead include the file contents.

If the input or output format is not specified, windres will guess based on the file name, or, for the input file, the file contents. A file with an extension of ‘.rc’ will be treated as an rc file, a file with an extension of ‘.res’ will be treated as a res file, and a file with an extension of ‘.o’ or ‘.exe’ will be treated as a coff file.

If no output file is specified, windres will print the resources in rc format to standard output.

The normal use is for you to write an rc file, use windres to convert it to a COFF object file, and then link the COFF file into your application. This will make the resources described in the rc file available to Windows.

-i filename

--input filename

The name of the input file. If this option is not used, then windres will use the first non-option argument as the input file name. If there are no non-option arguments, then windres will read from standard input. windres can not read a COFF file from standard input.

-o filename

--output filename

The name of the output file. If this option is not used, then windres will use the first non-option argument, after any used for the input file name, as the output file name. If there is no non-option argument, then windres will write to standard output. windres can not write a COFF file to standard output.

-I format

--input-format format

The input format to read. format may be ‘res’, ‘rc’, or ‘coff’. If no input format is specified, windres will guess, as described above.

-O format

--output-format format

The output format to generate. format may be ‘res’, ‘rc’, or ‘coff’. If no output format is specified, windres will guess, as described above.

-F target

--target target

Specify the BFD format to use for a COFF file as input or output. This is a BFD target name; you can use the --help option to see a list of supported targets. Normally windres will use the default format, which is the first one listed by the --help option. Section 16.1 [Target Selection], page 51.

--preprocessor program

When windres reads an rc file, it runs it through the C preprocessor first. This option may be used to specify the preprocessor to use, including any leading arguments. The default preprocessor argument is gcc -E -xc-header -DRC_

INVOKED.
--include-dir directory

Specify an include directory to use when reading an rc file. windres will pass this to the preprocessor as an -I option. windres will also search this directory when looking for files named in the rc file.

-D target

--define sym[=val]

Specify a -D option to pass to the preprocessor when reading an rc file.

-v Enable verbose mode. This tells you what the preprocessor is if you didn’t specify one.

--language val

Specify the default language to use when reading an rc file. val should be a hexadecimal language code. The low eight bits are the language, and the high eight bits are the sublanguage.

--use-temp-file

Use a temporary file to instead of using popen to read the output of the preprocessor. Use this option if the popen implementation is buggy on the host (eg., certain non-English language versions of Windows 95 and Windows 98 are known to have buggy popen where the output will instead go the console).

--no-use-temp-file

Use popen, not a temporary file, to read the output of the preprocessor. This is the default behaviour.

--help Prints a usage summary.

--version

Prints the version number for windres.

--yydebug

If windres is compiled with YYDEBUG defined as 1, this will turn on parser debugging.

44 gnu Binary Utilities

14 Create files needed to build and use DLLs

dlltool may be used to create the files needed to build and use dynamic link libraries (DLLs).

Warning: dlltool is not always built as part of the binary utilities, since it is only useful for those targets which support DLLs.

     dlltool [-d|--input-def def-file-name]
             [-b|--base-file base-file-name]
             [-e|--output-exp exports-file-name]
             [-z|--output-def def-file-name]
             [-l|--output-lib library-file-name]
             [--export-all-symbols] [--no-export-all-symbols]
             [--exclude-symbols list]
             [--no-default-excludes]
             [-S|--as path-to-assembler] [-f|--as-flags options]
             [-D|--dllname name] [-m|--machine machine]
             [-a|--add-indirect] [-U|--add-underscore] [-k|--kill-at]
             [-A|--add-stdcall-alias]
             [-x|--no-idata4] [-c|--no-idata5] [-i|--interwork]
             [-n|--nodelete] [-v|--verbose] [-h|--help] [-V|--version]
             [object-file ...]

dlltool reads its inputs, which can come from the ‘-d’ and ‘-b’ options as well as object files specified on the command line. It then processes these inputs and if the ‘-e’ option has been specified it creates a exports file. If the ‘-l’ option has been specified it creates a library file and if the ‘-z’ option has been specified it creates a def file. Any or all of the -e, -l and -z options can be present in one invocation of dlltool.

When creating a DLL, along with the source for the DLL, it is necessary to have three other files. dlltool can help with the creation of these files.

The first file is a ‘.def’ file which specifies which functions are exported from the DLL, which functions the DLL imports, and so on. This is a text file and can be created by hand, or dlltool can be used to create it using the ‘-z’ option. In this case dlltool will scan the object files specified on its command line looking for those functions which have been specially marked as being exported and put entries for them in the .def file it creates.

In order to mark a function as being exported from a DLL, it needs to have an ‘-export:<name_of_function>’ entry in the ‘.drectve’ section of the object file. This can be done in C by using the asm() operator:

       asm (".section .drectve");
       asm (".ascii \"-export:my_func\"");
       int my_func (void) { ... }

The second file needed for DLL creation is an exports file. This file is linked with the object files that make up the body of the DLL and it handles the interface between the DLL and the outside world. This is a binary file and it can be created by giving the ‘-e’ option to dlltool when it is creating or reading in a .def file.

The third file needed for DLL creation is the library file that programs will link with in order to access the functions in the DLL. This file can be created by giving the ‘-l’ option to dlltool when it is creating or reading in a .def file.

dlltool builds the library file by hand, but it builds the exports file by creating temporary files containing assembler statements and then assembling these. The ‘-S’ command line option can be used to specify the path to the assembler that dlltool will use, and the ‘-f’ option can be used to pass specific flags to that assembler. The ‘-n’ can be used to prevent dlltool from deleting these temporary assembler files when it is done, and if ‘-n’ is specified twice then this will prevent dlltool from deleting the temporary object files it used to build the library.

Here is an example of creating a DLL from a source file ‘dll.c’ and also creating a program (from an object file called ‘program.o’) that uses that DLL:

       gcc -c dll.c
       dlltool -e exports.o -l dll.lib dll.o
       gcc dll.o exports.o -o dll.dll
       gcc program.o dll.lib -o program

The command line options have the following meanings:

-d filename

--input-def filename

Specifies the name of a .def file to be read in and processed.

-b filename

--base-file filename

Specifies the name of a base file to be read in and processed. The contents of this file will be added to the relocation section in the exports file generated by dlltool.

-e filename

--output-exp filename

Specifies the name of the export file to be created by dlltool.

-z filename

--output-def filename

Specifies the name of the .def file to be created by dlltool.

-l filename

--output-lib filename

Specifies the name of the library file to be created by dlltool.

--export-all-symbols

Treat all global and weak defined symbols found in the input object files as symbols to be exported. There is a small list of symbols which are not exported by default; see the --no-default-excludes option. You may add to the list of symbols to not export by using the --exclude-symbols option.

--no-export-all-symbols

Only export symbols explicitly listed in an input .def file or in ‘.drectve’ sections in the input object files. This is the default behaviour. The ‘.drectve’ sections are created by ‘dllexport’ attributes in the source code.

--exclude-symbols list

Do not export the symbols in list. This is a list of symbol names separated by comma or colon characters. The symbol names should not contain a leading underscore. This is only meaningful when --export-all-symbols is used.

--no-default-excludes

When --export-all-symbols is used, it will by default avoid exporting certain special symbols. The current list of symbols to avoid exporting is ‘DllMain@12’,

‘DllEntryPoint@0’, ‘impure_ptr’. You may use the --no-default-excludes

option to go ahead and export these special symbols. This is only meaningful

when --export-all-symbols is used.

-S path --as path Specifies the path, including the filename, of the assembler to be used to create

the exports file.

-f switches

--as-flags switches

Specifies any specific command line switches to be passed to the assembler when building the exports file. This option will work even if the ‘-S’ option is not used. This option only takes one argument, and if it occurs more than once on the command line, then later occurrences will override earlier occurrences. So if it is necessary to pass multiple switches to the assembler they should be enclosed in double quotes.

-D name

--dll-name name

Specifies the name to be stored in the .def file as the name of the DLL when the ‘-e’ option is used. If this option is not present, then the filename given to the ‘-e’ option will be used as the name of the DLL.

-m machine

-machine machine

Specifies the type of machine for which the library file should be built. dlltool has a built in default type, depending upon how it was created, but this option can be used to override that. This is normally only useful when creating DLLs for an ARM processor, when the contents of the DLL are actually encode using THUMB instructions.

-a
--add-indirect

Specifies that when dlltool is creating the exports file it should add a section which allows the exported functions to be referenced without using the import library. Whatever the hell that means!

-U
--add-underscore

Specifies that when dlltool is creating the exports file it should prepend an underscore to the names of the exported functions.

-k
--kill-at

Specifies that when dlltool is creating the exports file it should not append the string ‘@ <number>’. These numbers are called ordinal numbers and they represent another way of accessing the function in a DLL, other than by name.

-A
--add-stdcall-alias

Specifies that when dlltool is creating the exports file it should add aliases for stdcall symbols without ‘@ <number>’ in addition to the symbols with ‘@

<number>’.
-x
--no-idata4

Specifies that when dlltool is creating the exports and library files it should omit the .idata4 section. This is for compatibility with certain operating systems.

-c
--no-idata5

Specifies that when dlltool is creating the exports and library files it should omit the .idata5 section. This is for compatibility with certain operating systems.

-i
--interwork

Specifies that dlltool should mark the objects in the library file and exports file that it produces as supporting interworking between ARM and THUMB code.

-n
--nodelete

Makes dlltool preserve the temporary assembler files it used to create the exports file. If this option is repeated then dlltool will also preserve the temporary object files it uses to create the library file.

-v
--verbose

Make dlltool describe what it is doing.

-h

--help Displays a list of command line options and then exits.

-V
--version

Displays dlltool’s version number and then exits.

15 readelf

     readelf [ -a | --all ]
             [ -h | --file-header]
             [ -l | --program-headers | --segments]
             [ -S | --section-headers | --sections]
             [ -e | --headers]
             [ -s | --syms | --symbols]
             [ -n | --notes]
             [ -r | --relocs]
             [ -d | --dynamic]
             [ -V | --version-info]
             [ -D | --use-dynamic]
             [ -x <number> | --hex-dump=<number>]
             [ -w[liapr] | --debug-dump[=info,=line,=abbrev,=pubnames,=ranges]]
             [      --histogram]
             [ -v | --version]
             [ -H | --help]

elffile...

readelf displays information about one or more ELF format object files. The options control what particular information to display.

elffile… are the object files to be examined. At the moment, readelf does not support examining archives, nor does it support examing 64 bit ELF files.

The long and short forms of options, shown here as alternatives, are equivalent. At least one option besides ‘-v’ or ‘-H’ must be given.

-a
--all  Equivalent to specifiying ‘--file-header’, ‘--program-headers’,
          ‘--sections’, ‘--symbols’, ‘--relocs’, ‘--dynamic’, ‘--notes’ and
          ‘--version-info’.
-h
--file-header

Displays the information contained in the ELF header at the start of the file.

-l
--program-headers
--segments

Displays the information contained in the file’s segment headers, if it has any.

-S
--sections
--section-headers

Displays the information contained in the file’s section headers, if it has any.

-s
--symbols

--syms Displays the entries in symbol table section of the file, if it has one.

-e
--headers

Display all the headers in the file. Equivalent to ‘-h -l -S’.

-n

--notes Displays the contents of the NOTE segment, if it exists.

-r

--relocs Displays the contents of the file’s relocation section, if it ha one.

-d
--dynamic

Displays the contents of the file’s dynamic section, if it has one.

-V
--version-info

Displays the contents of the version sections in the file, it they exist.

-D
--use-dynamic

When displaying symbols, this option makes readelf use the symbol table in the file’s dynamic section, rather than the one in the symbols section.

-x <number>
--hex-dump=<number>

Displays the contents of the indicated section as a hexadecimal dump.

-w[liapr]
--debug-dump[=line,=info,=abbrev,=pubnames,=ranges]

Displays the contents of the debug sections in the file, if any are present. If one of the optional letters or words follows the switch then only data found in those specific sections will be dumped.

--histogram

Display a histogram of bucket list lengths when displaying the contents of the symbol tables.

-v
--version

Display the version number of readelf.

-H

--help Display the command line options understood by readelf.

16 Selecting the target system

You can specify three aspects of the target system to the gnu binary file utilities, each in several ways:

• the target

• the architecture

• the linker emulation (which applies to the linker only)

In the following summaries, the lists of ways to specify values are in order of decreasing precedence. The ways listed first override those listed later.

The commands to list valid values only list the values for which the programs you are running were configured. If they were configured with ‘--enable-targets=all’, the commands list most of the available values, but a few are left out; not all targets can be configured in at once because some of them can only be configured native (on hosts with the same type as the target system).

16.1 Target Selection

A target is an object file format. A given target may be supported for multiple architectures (see Section 16.2 [Architecture Selection], page 52). A target selection may also have variations for different operating systems or architectures.

The command to list valid target values is ‘objdump -i’ (the first column of output contains the relevant information).

   Some sample values are: ‘a.out-hp300bsd’, ‘ecoff-littlemips’, ‘a.out-sunos-big’.

You can also specify a target using a configuration triplet. This is the same sort of name that is passed to ‘configure’ to specify a target. When you use a configuration triplet as an argument, it must be fully canonicalized. You can see the canonical version of a triplet by running the shell script ‘config.sub’ which is included with the sources.

Some sample configuration triplets are: ‘m68k-hp-bsd’, ‘mips-dec-ultrix’,

‘sparc-sun-sunos’.
objdump Target

Ways to specify:

1. command line option: ‘-b’ or ‘--target’

2. environment variable GNUTARGET

3. deduced from the input file

objcopy and strip Input Target

Ways to specify:

1. command line options: ‘-I’ or ‘--input-target’, or ‘-F’ or ‘--target’

2. environment variable GNUTARGET

3. deduced from the input file

objcopy and strip Output Target

Ways to specify:

1. command line options: ‘-O’ or ‘--output-target’, or ‘-F’ or ‘--target’

2. the input target (see “objcopy and strip Input Target” above)

3. environment variable GNUTARGET

4. deduced from the input file

nm, size, and strings Target

Ways to specify:

1. command line option: ‘--target’

2. environment variable GNUTARGET

3. deduced from the input file

Linker Input Target

Ways to specify:

1. command line option: ‘-b’ or ‘--format’ (see section “Options” in Using LD)

2. script command TARGET (see section “Option Commands” in Using LD)

3. environment variable GNUTARGET (see section “Environment” in Using LD)

4. the default target of the selected linker emulation (see Section 16.3 [Linker Emulation Selection], page 53)

Linker Output Target

Ways to specify:

1. command line option: ‘-oformat’ (see section “Options” in Using LD)

2. script command OUTPUT_FORMAT (see section “Option Commands” in Using LD)

3. the linker input target (see “Linker Input Target” above)

16.2 Architecture selection

An architecture is a type of cpu on which an object file is to run. Its name may contain a colon, separating the name of the processor family from the name of the particular cpu.

The command to list valid architecture values is ‘objdump -i’ (the second column contains the relevant information).

  Sample values: ‘m68k:68020’, ‘mips:3000’, ‘sparc’.

objdump Architecture

Ways to specify:

1. command line option: ‘-m’ or ‘--architecture’

2. deduced from the input file

objcopy, nm, size, strings Architecture

Ways to specify:

1. deduced from the input file

Linker Input Architecture

Ways to specify:

1. deduced from the input file

Linker Output Architecture

Ways to specify:

1. script command OUTPUT_ARCH (see section “Option Commands” in Using LD)

2. the default architecture from the linker output target (see Section 16.1 [Target Selection], page 51)

16.3 Linker emulation selection

A linker emulation is a “personality” of the linker, which gives the linker default values for the other aspects of the target system. In particular, it consists of

• the linker script

• the target

• several “hook” functions that are run at certain stages of the linking process to do special things that some targets require The command to list valid linker emulation values is ‘ld -V’. Sample values: ‘hp300bsd’, ‘mipslit’, ‘sun4’. Ways to specify:

1. command line option: ‘-m’ (see section “Options” in Using LD)

2. environment variable LDEMULATION

3. compiled-in DEFAULT_EMULATION from ‘Makefile’, which comes from EMUL in

 ‘config/target.mt’

54 gnu Binary Utilities

17 Reporting Bugs

Your bug reports play an essential role in making the binary utilities reliable.

Reporting a bug may help you by bringing a solution to your problem, or it may not. But in any case the principal function of a bug report is to help the entire community by making the next version of the binary utilities work better. Bug reports are your contribution to their maintenance.

In order for a bug report to serve its purpose, you must include the information that enables us to fix the bug.

17.1 Have you found a bug?

If you are not sure whether you have found a bug, here are some guidelines:

• If a binary utility gets a fatal signal, for any input whatever, that is a bug. Reliable utilities never crash.

• If a binary utility produces an error message for valid input, that is a bug.

• If you are an experienced user of binary utilities, your suggestions for improvement are welcome in any case.

17.2 How to report bugs

A number of companies and individuals offer support for gnu products. If you obtained the binary utilities from a support organization, we recommend you contact that organization first.

You can find contact information for many support companies and individuals in the file ‘etc/SERVICE’ in the gnu Emacs distribution.

In any event, we also recommend that you send bug reports for the binary utilities to

‘bug-gnu-utils@gnu.org’.

The fundamental principle of reporting bugs usefully is this: report all the facts. If you are not sure whether to state a fact or leave it out, state it!

Often people omit facts because they think they know what causes the problem and assume that some details do not matter. Thus, you might assume that the name of a file you use in an example does not matter. Well, probably it does not, but one cannot be sure. Perhaps the bug is a stray memory reference which happens to fetch from the location where that pathname is stored in memory; perhaps, if the pathname were different, the contents of that location would fool the utility into doing the right thing despite the bug. Play it safe and give a specific, complete example. That is the easiest thing for you to do, and the most helpful.

Keep in mind that the purpose of a bug report is to enable us to fix the bug if it is new to us. Therefore, always write your bug reports on the assumption that the bug has not been reported previously.

Sometimes people give a few sketchy facts and ask, “Does this ring a bell?” Those bug reports are useless, and we urge everyone to refuse to respond to them except to chide the sender to report bugs properly.

To enable us to fix the bug, you should include all these things:

• The version of the utility. Each utility announces it if you start it with the ‘--version’ argument. Without this, we will not know whether there is any point in looking for the bug in the current version of the binary utilities.

• Any patches you may have applied to the source, including any patches made to the BFD library.

• The type of machine you are using, and the operating system name and version number.

• What compiler (and its version) was used to compile the utilities—e.g. “gcc-2.7”.

• The command arguments you gave the utility to observe the bug. To guarantee you will not omit something important, list them all. A copy of the Makefile (or the output from make) is sufficient. If we were to try to guess the arguments, we would probably guess wrong and then we might not encounter the bug.

• A complete input file, or set of input files, that will reproduce the bug. If the utility is reading an object file or files, then it is generally most helpful to send the actual object files, uuencoded if necessary to get them through the mail system. Note that ‘bug-gnu-utils@gnu.org’ is a mailing list, so you should avoid sending very large files to it. Making the files available for anonymous FTP is OK. If the source files were produced exclusively using gnu programs (e.g., gcc, gas, and/or the gnu ld), then it may be OK to send the source files rather than the object files. In this case, be sure to say exactly what version of gcc, or whatever, was used to produce the object files. Also say how gcc, or whatever, was configured.

• A description of what behavior you observe that you believe is incorrect. For example, “It gets a fatal signal.” Of course, if the bug is that the utility gets a fatal signal, then we will certainly notice it. But if the bug is incorrect output, we might not notice unless it is glaringly wrong. You might as well not give us a chance to make a mistake. Even if the problem you experience is a fatal signal, you should still say so explicitly. Suppose something strange is going on, such as your copy of the utility is out of synch, or you have encountered a bug in the C library on your system. (This has happened!) Your copy might crash and ours would not. If you told us to expect a crash, then when ours fails to crash, we would know that the bug was not happening for us. If you had not told us to expect a crash, then we would not be able to draw any conclusion from our observations.

• If you wish to suggest changes to the source, send us context diffs, as generated by diff with the ‘-u’, ‘-c’, or ‘-p’ option. Always send diffs from the old file to the new file. If you wish to discuss something in the ld source, refer to it by context, not by line number. The line numbers in our development sources will not match those in your sources. Your line numbers would convey no useful information to us.

Here are some things that are not necessary:

• A description of the envelope of the bug.

Often people who encounter a bug spend a lot of time investigating which changes to the input file will make the bug go away and which changes will not affect it.

This is often time consuming and not very useful, because the way we will find the bug is by running a single example under the debugger with breakpoints, not by pure deduction from a series of examples. We recommend that you save your time for something else. Of course, if you can find a simpler example to report instead of the original one, that is a convenience for us. Errors in the output will be easier to spot, running under the debugger will take less time, and so on. However, simplification is not vital; if you do not want to do this, report the bug anyway and send us the entire test case you used.

• A patch for the bug.

A patch for the bug does help us if it is a good one. But do not omit the necessary information, such as the test case, on the assumption that a patch is all we need. We might see problems with your patch and decide to fix the problem another way, or we might not understand it at all. Sometimes with programs as complicated as the binary utilities it is very hard to construct an example that will make the program follow a certain path through the code. If you do not send us the example, we will not be able to construct one, so we will not be able to verify that the bug is fixed. And if we cannot understand what bug you are trying to fix, or why your patch should be an improvement, we will not install it. A test case will help us to understand.

• A guess about what the bug is or what it depends on. Such guesses are usually wrong. Even we cannot guess right about such things without first using the debugger to find the facts.

58 gnu Binary Utilities Index 59

Index

.

.stab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

A

addr2line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 address to file name and line number . . . . . . . . . . 37 all header information, object file . . . . . . . . . . . . . 25 ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ar compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 architectures available. . . . . . . . . . . . . . . . . . . . . . . . 23 archive contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 archive headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

B

base files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 bug criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 bug reports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

C

c++filt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 changing object addresses . . . . . . . . . . . . . . . . . . . . 18 changing section address . . . . . . . . . . . . . . . . . . . . . 18 changing section LMA . . . . . . . . . . . . . . . . . . . . . . . 18 changing section VMA . . . . . . . . . . . . . . . . . . . . . . . 18 changing start address . . . . . . . . . . . . . . . . . . . . . . . 18 collections of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 compatibility, ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 contents of archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 crash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 creating archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 cxxfilt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

D

dates in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 debug symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 debugging symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 deleting from archive . . . . . . . . . . . . . . . . . . . . . . . . . . 4 demangling C++ symbols . . . . . . . . . . . . . . . . . . . . . 35 demangling in nm . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 demangling in objdump . . . . . . . . . . . . . . . . . . . 22, 37 disassembling object code . . . . . . . . . . . . . . . . . . . . 22 disassembly architecture. . . . . . . . . . . . . . . . . . . . . . 23 disassembly endianness . . . . . . . . . . . . . . . . . . . . . . . 23 disassembly, with source. . . . . . . . . . . . . . . . . . . . . . 24 discarding symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 DLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 dlltool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 dynamic relocation entries, in object file . . . . . . . 24 dynamic symbol table entries, printing . . . . . . . . 25 dynamic symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

E

ELF core notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 ELF dynamic section information . . . . . . . . . . . . . 50 ELF file header information . . . . . . . . . . . . . . . . . . 49 ELF file information . . . . . . . . . . . . . . . . . . . . . . . . . 49 ELF object file format . . . . . . . . . . . . . . . . . . . . . . . 25 ELF program header information . . . . . . . . . . . . . 49 ELF reloc information . . . . . . . . . . . . . . . . . . . . . . . 50 ELF section information. . . . . . . . . . . . . . . . . . . . . . 49 ELF segment information . . . . . . . . . . . . . . . . . . . . 49 ELF symbol table information . . . . . . . . . . . . . . . . 49 ELF version sections informations. . . . . . . . . . . . . 50 endianness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 error on valid input . . . . . . . . . . . . . . . . . . . . . . . . . . 55 external symbols . . . . . . . . . . . . . . . . . . . . . . 12, 13, 14 extract from archive . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

fatal signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

H

header information, all . . . . . . . . . . . . . . . . . . . . . . . 25

I

input .def file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 input file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

L

ld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 linker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 listings strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

M

machine instructions . . . . . . . . . . . . . . . . . . . . . . . . . 22 moving in archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 MRI compatibility, ar . . . . . . . . . . . . . . . . . . . . . . . . . 6

60 gnu Binary Utilities

N

name duplication in archive. . . . . . . . . . . . . . . . . . . . 5 name length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 nm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 nm compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 nm format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 not writing archive index . . . . . . . . . . . . . . . . . . . . . . 6

O

objdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 object code format . . . . . . . . . . . . . 13, 22, 30, 31, 37 object file header . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 object file information. . . . . . . . . . . . . . . . . . . . . . . . 21 object file sections . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 object formats available . . . . . . . . . . . . . . . . . . . . . . 23 operations on archive . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

printing from archive. . . . . . . . . . . . . . . . . . . . . . . . . . 4 printing strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

Q

quick append to archive . . . . . . . . . . . . . . . . . . . . . . . 4

R

radix for section sizes . . . . . . . . . . . . . . . . . . . . . . . . 30 ranlib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 readelf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 relative placement in archive. . . . . . . . . . . . . . . . . . . 5 relocation entries, in object file . . . . . . . . . . . . . . . 24 removing symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 repeated names in archive . . . . . . . . . . . . . . . . . . . . . 5 replacement in archive . . . . . . . . . . . . . . . . . . . . . . . . 4 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

S

scripts, ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 section addresses in objdump . . . . . . . . . . . . . . . . . 22 section headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 section information . . . . . . . . . . . . . . . . . . . . . . . . . . 23 section sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 sections, full contents . . . . . . . . . . . . . . . . . . . . . . . . 24 size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 size display format . . . . . . . . . . . . . . . . . . . . . . . . . . 29 size number format . . . . . . . . . . . . . . . . . . . . . . . . . 30 sorting symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 source code context . . . . . . . . . . . . . . . . . . . . . . . . . . 23 source disassembly . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 source file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 source filenames for object files . . . . . . . . . . . . . . . 23 stab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 start-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 stop-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 strings, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 symbol index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 27 symbol index, listing . . . . . . . . . . . . . . . . . . . . . . . . . 13 symbol line numbers . . . . . . . . . . . . . . . . . . . . . . . . . 13 symbol table entries, printing . . . . . . . . . . . . . . . . . 25 symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 symbols, discarding . . . . . . . . . . . . . . . . . . . . . . . . . . 33

U

undefined symbols . . . . . . . . . . . . . . . . . . . . . . . . 13, 14 Unix compatibility, ar. . . . . . . . . . . . . . . . . . . . . . . . . 4 updating an archive . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

V

version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 VMA in objdump. . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

W

wide output, printing . . . . . . . . . . . . . . . . . . . . . . . . 25 writing archive index . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 ar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Controlling ar on the command line . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Controlling ar with a script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 ld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3 nm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4 objcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5 objdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6 ranlib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7 size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 8 strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9 strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 10 c++filt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 11 addr2line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 12 nlmconv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 13 windres . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 14 Create files needed to build and use DLLs

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 15 readelf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 16 Selecting the target system. . . . . . . . . . . . . . . 51

16.1 Target Selection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 16.2 Architecture selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 16.3 Linker emulation selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

17 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . 55

17.1 Have you found a bug? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 17.2 How to report bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59